a 10 kg disk (I+1/2R^2) starts to roll w/o slipping down a 5m tall incline. what is its liner velocity at the bottom of the incline?

answer = 8.1 m/s

pls show work and steps
Answer #1

inital energy Ei = m*g*h

final energy energy at the bottom Ef = KEt + KEr


KEt = translational kinetic energy = (1/2)*m*v^2


KEr = rotatioanl kietic energy = (1/2)*I*w^2

I = momet of inertia= (1/2)*m*r^2


w = angular speed = v/r


therefore

Ef = (1/2)*m*v^2 + (1/2)*(1/2)*m*r^2*(v/r)^2


Ef = (1/2)*m*v^2 + (1/4)*m*v^2


Ef = (3/4)*m*v^2


from energy conservation


total energy is constant


Ef = Ei


(3/4)*m*v^2 = m*g*h


v = sqrt(4gh/3)


v = sqrt(4*9.81*5/3)


v = 8.1 m/s